Need help to resolve. Thanks please!!
6
∑ (n+1)
n=2

Answer by fractalier(6550)

Okay, the summation of (n + 1) as n goes from 2 to 6 becomes
(2+1) + (3+1) + (4+1) + (5+1) + (6+1) =
3 + 4 + 5 + 6 + 7 =
25